Sir,
Pl. help me to solve this problem
Length of a rectangular field is 23 mtrs more than its Width. If its Area is 420 sq. mtrs, find its length and width.
I tried like this
Let Width be = x mtrs
Then Length is = x+23 mtrs
(x)+ (x+23) = 420 mtrs
x2 + 23x = 420 mtrs
I could not find the value of x. Pl. help me.
Regards

Answer by praseenakos@yahoo.com(507)

area of the rectangle is = 420
X^2+23X = 420
X^2+23X-420 = 0 by factorising 420 we get (35).(-12)
therefore X^2+35X-12X-420=0
X(X+35)-12(X+35)=0 or (X+35).(X-12)=0 X-12=0
X=12
X+35 =0 X=-35
the solution can be written as follows
1 if X=12 then Y= 12+23=35
2 if X=-35 then Y=12-35=-23